# Fairness in Ranking under Disparate Uncertainty

### v3

@@ E.2. Synthetic Dataset
  • Low Disparate Uncertainty: $Beta(\frac{1}{20},\frac{1}{20})$. Note that even though both groups are drawn from the same distribution, any sampled instance still contains some amount of disparate uncertainty.
@@
We compute the costs using $\mathbb{P}(r_{i}|\mathcal{D}),\mathbb{P}(i\in\sigma_{k}^{\pi})$ according to Equation 3, and 4.
Next, we plot a random sample from Table 1 according to the generation process described above in Figure 11. We also qualitatively analyze a commonly used measure of utility to the principal, namely, the expected Normalized Discounted Cumulative Gain (nDCG), which according to our model is,
@@
As shown in Figure 10, the nDCG for EOR is only slightly lower than the nDCG optimal PRP ranking and competitive with all other ranking policies.
@@ E.3. US Census Survey Dataset
Note the overlapping of $\pi^{RA}$ and $\pi^{PRP}$ in Figure 13 and 14. This is expected because $\pi^{RA}$ swaps the candidates in PRP ranking to satisfy proportional exposure as described in Section E.1. For USCensus dataset, amortized exposure is already satisfied between groups, but this does not imply that candidates are selected in proportion to their expected relevance.

### v4

@@ E.2. Synthetic Dataset
  • Low Disparate Uncertainty: $Beta(\frac{1}{20},\frac{1}{20})$. Note that even when both groups are drawn from the same distribution, any sampled instance still contains some amount of disparate uncertainty.
@@
We compute the costs using $\mathbb{P}(r_{i}|\mathcal{D}),\mathbb{P}(i\in\sigma_{k}^{\pi})$ according to Eqs. (3), and (4).
In Figure 11, we plot a random sample from Table 1 according to the generation process described above. We also qualitatively analyze a commonly used measure of utility to the principal, namely, the expected Normalized Discounted Cumulative Gain (nDCG), which according to our model is,
@@
As shown in Figure 10, the nDCG for EOR ranking is only slightly lower than the nDCG optimal PRP ranking and competitive with all other ranking policies.
@@ E.3. US Census Survey Dataset
Note the overlapping of $\pi^{RA}$ and $\pi^{PRP}$ in Figure 13 and 14. This is expected because $\pi^{RA}$ swaps the candidates in PRP ranking to satisfy proportional exposure as described in Appendix E.1. Since the amortized exposure between groups is already satisfied with the PRP ranking for this dataset, $\pi^{RA}$ and $\pi^{PRP}$ compute similar rankings.